# The Ordered Zeckendorf Game

### v1

@@ 1.1. History and Motivation
The Fibonacci sequence is given by $F_{1}=1,F_{2}=2$, and $F_{i}=F_{i}+F_{i-1}$ for $i\geq 2$. Zeckendorf’s Theorem [9] states that every positive integer can be uniquely represented as a sum of non-consecutive Fibonacci numbers, noting that in order to guarantee uniqueness it is necessary that we have only one term equal to $1$. This decomposition can be efficiently obtained via a greedy algorithm, and its structural and probabilistic properties have been extensively investigated across combinatorics, number theory, and discrete mathematics. A game-theoretic interpretation of the theorem was first proposed in [1], where two players alternately perform legal replacements on a multiset of Fibonacci numbers until reaching the unique Zeckendorf decomposition. A formal definition of the classical Zeckendorf Game is given below.
@@ 1.2. Main Results
In particular, there is at most one repetition in the higher-index terms. A state-transition graph is shown in Figure 1.
@@
Figure 1. Directed graph showing the possible ways to move between game states in Lemma 1.9
@@
$M(n)~\sim~\frac{n^{2}}{2}\quad\text{as }n\to\infty.$
Moreover, we can estimate an error by
$\dfrac{n^{2}}{2}-M(n)~=~\Theta(n\log_{\phi}(n)).$
@@ 3. Game Length: More Bounds and Extremes
In the first case, we remain in the same game state. In the second, if there is no $F_{3}$ at the beginning of higher-index terms, then we still remain in No Repetitions, otherwise we transition into State 1.
@@
For the lower bound, we begin the game by merging the leftmost two copies of $F_{1}$, then moving the copy of $F_{2}$ created by this move to the rightmost end of the tuple via n2n-2 switch moves. We are left with the tuple $(F_{1},F_{1},...,F_{1},F_{2})$. As no more switch moves remain, we repeat this process: letting $k$ denote the number of copies of $F_{1}$ remaining, we merge the leftmost two copies of $F_{1}$ into a copy of $F_{2}$ in a single move and then perform k2k-2 switch moves to move the new copy of $F_{2}$ to the right of the k2k-2 remaining copies of $F_{1}$, yielding in k1k-1 moves a tuple consisting of k2k-2 copies of $F_{1}$ to the left of $\frac{n-k-2}{2}$ copies of $F_{2}$. We continue iterating this process until $k\leq 2$, noting that we begin with $k=n$ and decrement $k$ by $2$ at each iteration. Thus, this portion of the games terminates in $\sum_{j=0}^{\lfloor\frac{n-2}{2}\rfloor}n-2j-1$ moves. If $n$ is odd, this summation is equal to $(n-1)+(n-3)+...+1=\frac{n^{2}}{4}$ moves and this portion of the game terminates at the tuple $(F_{1},F_{2},F_{2},...F_{2})$, and if $n$ is even, this summation is equal to $(n-1)+(n-3)+...+2=\frac{(n-1)^{2}}{4}+\frac{n-1}{2}=\frac{n^{2}-1}{4}$ and this portion of the game terminates at the tuple $(F_{2},F_{2},...,F_{2})$.
@@
Now, we suppose the tuple has size $k$ and we write it as $(F_{2},\dots,F_{2},\textit{higher-index terms})$, where higher-index terms denote terms that are at least $F_{3}$. We consider the number of moves taken to convert it into the (k1)(k-1)-tuple $(F_{2},\dots,F_{2},\textit{higher-index terms})$. Note that the first $k$-tuple only contains a copy of $F_{1}$ when $n$ is odd and $k=\lceil\frac{n}{2}\rceil$. We ignore this case in our calculations as it only increases game length. By Lemma 1.9, there is at most one repeated higher-index term, so the number of higher-index terms is bounded by $\ell-1$, where $F_{\ell}$ is the maximal Fibonacci summand in the Zeckendorf decomposition of $n$. By [1], $\ell\leq\log_{\phi}(\sqrt{5}n+1/2)$, so the number of higher-index terms is bounded by
$c_{n}~\coloneq~\log_{\phi}(\sqrt{5}n+1/2)-1.$ (1)
Hence, there are at least $k-c_{n}-1$ lots of $F_{2}$ in the $k$-tuple.
@@
We assume that there are at least four lots of $F_{2}$ and that there are no lots of $F_{1}$. This holds for $c_{n}+5\leq k\leq\lfloor\frac{n}{2}\rfloor$. Then after reordering and splitting the higher-index terms, we have two cases:
  1. (i)
    $F_{3}$ is repeated,
  2. (ii)
    no higher-index terms are repeated.
Tables 1 and 2 show the paths to the state where the tuple has a copy of $F_{1}$ on the left in each of cases (i) and (ii).
State Moves
$(F_{2},\dots,F_{2},F_{3},F_{3},\textit{higher-index terms})$ 0
$(F_{2},\dots,F_{2},F_{1},F_{4},\textit{higher-index terms})$ 1
$(F_{1},F_{2},\dots,F_{2},F_{4},\textit{higher-index terms})$ $\geq k-c_{n}$
@@
Table 1. Sequence of moves to introduce $F_{1}$ on the left in case (ii).
State Moves
@@
$(F_{2},\dots,F_{2},F_{2},F_{2},\textit{higher-index terms})$ 0
@@
$(F_{2},\dots,F_{2},F_{1},F_{3},\textit{higher-index terms})$ 1
$(F_{1},F_{2},\dots,F_{2},F_{3},\textit{higher-index terms})$ $\geq k-c_{n}-2$
Table 2. Sequence of moves to introduce $F_{1}$ on the left in case (ii).
After some reordering moves and merging of higher-index terms, we have one of the following cases:
  1. (a)
    $F_{3}$ is repeated,
  2. (b)
    no higher-index terms are repeated.
The sequences of moves in each case are shown in Tables 3 and 4.
State Moves
$(F_{1},F_{2},\dots,F_{2},F_{3},F_{3},\textit{higher-index terms})$ 0
$(F_{1},F_{2},\dots,F_{2},F_{1},F_{4},\textit{higher-index terms})$ 1
$(F_{1},F_{1},F_{2},\dots,F_{2},F_{4},\textit{higher-index terms})$ $\geq k-c_{n}-2$
$(F_{2},F_{2},\dots,F_{2},F_{4},\textit{higher-index terms})$ $\geq k-c_{n}-1$
Table 3. Sequence of moves until the first merge move in case (a).
State Moves
$(F_{1},F_{2},\dots,F_{2},F_{2},F_{2},\textit{higher-index terms})$ 0
$(F_{1},F_{2},\dots,F_{2},F_{1},F_{3},\textit{higher-index terms})$ 1
$(F_{1},F_{1},F_{2},\dots,F_{2},F_{3},\textit{higher-index terms})$ $\geq k-c_{n}-4$
$(F_{2},F_{2},\dots,F_{2},F_{3},\textit{higher-index terms})$ $\geq k-c_{n}-3$
Table 4. Sequence of moves until the first merge move in case (b).
This shows that the number of moves between the game state becoming a $k$-tuple and a k1k-1-tuple is at least $2k-2c_{n}-5$, with $c_{n}$ given by (1).
Now we count the number of moves:
$M(n)$ $~\geq~\frac{n^{2}}{4}+\sum_{k=c_{n}+5}^{\lfloor\frac{n}{2}\rfloor}(2k-2c_{n}-5)$
$~=~\frac{n^{2}}{4}+\lfloor\frac{n}{2}\rfloor\left(\lfloor\frac{n}{2}\rfloor+1\right)-(c_{n}+5)(c_{n}+6)-(2c_{n}+5)\left(\lfloor\frac{n}{2}\rfloor-c_{n}-5\right)$
$~=~\frac{n^{2}}{2}-nc_{n}+o(n\log(n))$
@@
We also remark that in the proof above, we can improve a constant before $n\log_{\phi}(n)$ term up to $1/(8\phi)\approx 0.077$ by noting that, asymptotically, $F_{l}\geq n/\phi$ and $l\geq\log_{\phi}(n)$. However, assuming that LGS algorithm indeed gives a maximal game length, numerical simulations show that the constant before the $n\log_{\phi}(n)$ term is at least 0.4 and should be much lower that 1, as demonstrated on Figure 4. Hence, while Theorem 1.10 and Theorem 1.11 allow us to establish asymptotic behaviour of $\dfrac{n^{2}}{2}-M(n)$, they contain too crude estimates to give an insight for the explicit constant before $n\log_{\phi}(n)$.
@@
As $n\to\infty$, the distribution of the number of moves in a randomly played game, where all legal moves are equally likely, converges to a log-Gaussian distribution. Under these conditions, the game dynamics are sufficiently symmetric such that each player is equally likely to win.
This conjecture is supported by simulation data presented in Figure 2 and Figure 3, which suggest that the observed move counts exhibit the heavy-tailed, right-skewed characteristics consistent with a log-Gaussian distribution. Furthermore, the symmetry in the outcome frequencies aligns with the theoretical expectation of equal win probabilities for both players under a uniformly random move selection process.
@@ 5. Conclusion
  • Probabilistic Analysis of Random Play: Our empirical evidence suggests a log-Gaussian distribution of move counts under uniform random play. Can this distribution be derived rigorously? What are the deeper probabilistic and ergodic properties of the random dynamics?
@@ 6. Appendix
converges to a number bigger than 0.4 and much lower than 1.

### v2

@@ 1.1. History and Motivation
The Fibonacci sequence is given by $F_{1}=1,F_{2}=2$, and $F_{i+1}=F_{i}+F_{i-1}$ for $i\geq 2$. Zeckendorf’s Theorem [9] states that every positive integer can be uniquely represented as a sum of non-consecutive Fibonacci numbers, noting that in order to guarantee uniqueness it is necessary that we have only one term equal to $1$. This decomposition can be efficiently obtained via a greedy algorithm, and its structural and probabilistic properties have been extensively investigated across combinatorics, number theory, and discrete mathematics. A game-theoretic interpretation of the theorem was first proposed in [1], where two players alternately perform legal replacements on a multiset of Fibonacci numbers until reaching the unique Zeckendorf decomposition. A formal definition of the classical Zeckendorf Game is given below.
@@ 1.2. Main Results
In particular, there is at most one repetition in the higher-index terms. We display the possible ways to move between the states ”No Repetitions”, 1, 2, and 3 in a state-transition graph in Figure 1.
@@
Figure 1. Directed graph showing the possible ways to move between game states in Lemma 1.9.
@@
$M(n)~=~\frac{n^{2}}{2}-\Theta(n\log_{\phi}(n)).$
@@ 3. Game Length: More Bounds and Extremes
If there is no $F_{3}$ at the beginning of higher-index terms, then we still remain in No Repetitions, otherwise we transition into State 1.
@@
We begin the game by merging the leftmost two copies of $F_{1}$, then moving the copy of $F_{2}$ created by this move to the rightmost end of the tuple via n2n-2 switch moves. We are left with the tuple $(F_{1},F_{1},\dots,F_{1},F_{2})$. As no more switch moves remain, we repeat this process: letting $k$ denote the number of copies of $F_{1}$ remaining, we merge the leftmost two copies of $F_{1}$ into a copy of $F_{2}$ in a single move and then perform k2k-2 switch moves to move the new copy of $F_{2}$ to the right of the k2k-2 remaining copies of $F_{1}$, yielding in k1k-1 moves a tuple consisting of k2k-2 copies of $F_{1}$ to the left of $\frac{n-k-2}{2}$ copies of $F_{2}$. We continue iterating this process until $k\leq 2$, noting that we begin with $k=n$ and decrement $k$ by $2$ at each iteration. Thus, this portion of the games terminates in $\sum_{j=0}^{\lfloor\frac{n-2}{2}\rfloor}(n-2j-1)$ moves. If $n$ is odd, this summation is equal to $(n-1)+(n-3)+\dots+1=\frac{n^{2}}{4}$ moves and this portion of the game terminates at the tuple $(F_{1},F_{2},F_{2},\dots,F_{2})$, and if $n$ is even, this summation is equal to $(n-1)+(n-3)+\dots+2=\frac{(n-1)^{2}}{4}+\frac{n-1}{2}=\frac{n^{2}-1}{4}$ and this portion of the game terminates at the tuple $(F_{2},F_{2},\dots,F_{2})$.
@@
Now, we suppose the tuple has size $k$ and we write it as $(F_{2},\dots,F_{2},\textit{higher-index terms})$, where higher-index terms denotes higher-index terms that are at least $F_{3}$. We consider the number of moves taken to convert it into the (k1)(k-1)-tuple $(F_{2},\dots,F_{2},\textit{higher-index terms})$. Note that the first $k$-tuple only contains a copy of $F_{1}$ when $n$ is odd and $k=\lceil\frac{n}{2}\rceil$. We ignore this case in our calculations as it only increases game length. By Lemma 1.9, there is at most one repeated higher-index term, so the number of higher-index terms is bounded by $\ell-1$, where $F_{\ell}$ is the maximal Fibonacci summand in the Zeckendorf decomposition of $n$. By [2], $\ell\leq\log_{\phi}(\sqrt{5}n+1/2)$, so the number of higher-index terms is bounded by
$c_{n}:=\log_{\phi}(\sqrt{5}n+1/2)-1.$ (1)
Hence, there are at least $k-c_{n}-1$ copies of $F_{2}$ in the $k$-tuple.
@@
We assume that there are at least four copies of $F_{2}$ and that there are no copies of $F_{1}$. This holds for $c_{n}+5\leq k\leq\lfloor\frac{n}{2}\rfloor$.
[…]
Figure 2. State-transition graph of the ordered Zeckendorf Game with lower-order term tracking.
Edges: Full lines indicate constant tuple length; dotted lines indicate length decreases by one.
Nodes: Labels follow Lemma 1.9; green backgrounds indicate pivot states.
Subscripts: $xx_{1}$ denotes a lower-order tuple $(F_{1},F_{2},\dots,F_{2})$, while $xx$ denotes a lower-order tuple $(F_{2},\dots,F_{2})$.
Based on the bipartite state graph shown in Figure 2, we define two possible pivot states for this configuration, determined by the higher-index terms:
  • $F_{3}$ is repeated (Pivot State 1),
  • No higher-index terms are repeated (Pivot State NR).
Notice that along each path from a pivot state back to itself or to another pivot state, the length of the tuple decreases by exactly one. For simplicity, we use a lower bound of 0 moves for any transitions originating from states $2_{1},3_{1},2,\text{ and }3$. Furthermore, because the initial tuple consists entirely of $F_{2}$ as its lower-index terms, if the game begins in state $2$ or $3$, we assume an automatic transition to state $NR$.
Let us consider all possible paths between pivot states.
First, assume we begin at Pivot State 1. Table 1 illustrates the paths from Pivot State 1 to a state where the tuple has acquired a copy of $F_{1}$ on the far left.
@@
Table 1. Sequence of moves to introduce $F_{1}$ on the left from Pivot State 1.
Path State Moves
@@
Shared $(F_{2},\dots,F_{2},F_{3},F_{3},\textit{higher-index terms}_{\geq 4})$ (from Pivot State $1$) 0
Sequence $(F_{2},\dots,F_{2},F_{1},F_{4},\textit{higher-index terms}_{\geq 4})$ 1
Path A $(F_{1},F_{2},\dots,F_{2},F_{4},\textit{higher-index terms}_{\geq 5})$ (to $NR_{1}$) $\geq k-c_{n}-1$
Path B $(F_{1},F_{2},\dots,F_{2},F_{4},F_{4},\textit{higher-index terms}_{\geq 5})$ (to $2_{1}$) $\geq k-c_{n}-1$
$(F_{1},F_{2},\dots,F_{2},F_{2},F_{5},\textit{higher-index terms}_{\geq 5})$ (to $3_{1}$ or $NR_{1}$) $\geq k-c_{n}$
$(F_{1},F_{2},\dots,F_{2},F_{2},\textit{higher-index terms})$ (to $NR_{1}$) $\geq k-c_{n}$
Following these moves, the tuple reaches state $NR_{1}$ after a minimum of $k-c_{n}-1$ moves. By appending an $F_{1}$ to the left end, the total number of lower-index terms becomes at least $k-c_{n}$. From state $NR_{1}$, the sequence must subsequently return to a pivot state. Table 2 outlines these possible moves.
Table 2. Sequence of moves from $NR_{1}$ to a Pivot State.
Path State Moves
Shared $(F_{1},F_{2},\dots,F_{2},F_{2},\textit{higher-index terms})$ (from $NR_{1}$) 0
Path A $(F_{1},\dots,F_{2},F_{1},F_{3},\textit{higher-index terms}_{i\geq 4})$ 1
$(F_{1},F_{1},\dots,F_{2},F_{3},\textit{higher-index terms}_{i\geq 4})$ $\geq k-c_{n}-2$
$(F_{2},\dots,F_{2},F_{3},\textit{higher-index terms}_{i\geq 4})$ (to Pivot State NR) $\geq k-c_{n}-1$
Path B $(F_{1},\dots,F_{2},F_{1},F_{3},F_{3},\textit{higher-index terms}_{i\geq 4})$ 11
$(F_{1},F_{1},\dots,F_{3},F_{3},\textit{higher-index terms}_{i\geq 4})$ $\geq k-c_{n}-2$
$(F_{2},\dots,F_{2},F_{3},F_{3},\textit{higher-index terms}_{i\geq 4})$ (to Pivot State 1) $\geq k-c_{n}-1$
Consequently, transitioning from Pivot State 1 to another pivot state requires a minimum of $2k-2c_{n}-2$ moves. Completion of this path corresponds to a decrease of exactly 1 in the tuple’s length.
Next, suppose we begin at Pivot State NR. Table 3 displays the possible paths originating from this state.
Table 3. Sequence of moves from Pivot State NR to other states.
Path State Moves
@@
Shared $(F_{2},\dots,F_{2},F_{2},F_{2},\textit{higher-index terms})$ (from Pivot State NR) 0
Path A $(F_{2},\dots,F_{2},F_{1},F_{3},\textit{higher-index terms}_{\geq 4})$ 1
$(F_{1},F_{2},\dots,F_{2},F_{3},\textit{higher-index terms}_{\geq 4})$ (to $NR_{1}$) $\geq k-c_{n}-2$
Path B $(F_{2},F_{2},\dots,F_{1},F_{3},F_{3},\textit{higher-index terms}_{\geq 4})$ 11
$(F_{1},F_{2},\dots,F_{3},F_{3},\textit{higher-index terms}_{\geq 4})$ (to $1_{1}$) $\geq k-c_{n}-2$
If this sequence ends in state $NR_{1}$, the number of lower-index terms remains unchanged. Therefore, we can estimate the remaining moves required to reach the next pivot state by taking the lower bound from Table 2 and subtracting 1.
Alternatively, if the sequence ends in state $1_{1}$, the required moves to return to Pivot State NR are detailed in Table 4.
Table 4. Sequence of moves from $1_{1}$ to Pivot State NR.
Path State Moves
Shared $(F_{1},F_{2},\dots,F_{2},F_{3},F_{3},\textit{higher-index terms})$ (from $1_{1}$) 0
Path A $(F_{1},F_{2},\dots,F_{2},F_{1},F_{4},\textit{higher-index terms}_{\geq 5})$ 1
$(F_{1},F_{1},F_{2},\dots,F_{2},F_{4},\textit{higher-index terms}_{\geq 5})$ $\geq k-c_{n}-2$
$(F_{2},F_{2},\dots,F_{2},F_{4},\textit{higher-index terms}_{\geq 5})$ (to Pivot State NR) $\geq k-c_{n}-1$
Path B $(F_{1},F_{2},\dots,F_{2},F_{1},F_{4},F_{4},\textit{higher-index terms}_{\geq 5})$ 1
$(F_{1},F_{1},F_{2},\dots,F_{4},F_{4},\textit{higher-index terms}_{\geq 5})$ $\geq k-c_{n}-2$
$(F_{2},F_{2},\dots,F_{4},F_{4},\textit{higher-index terms}_{\geq 5})$ (to $2$) $\geq k-c_{n}-1$
$(F_{2},F_{2},\dots,F_{2},F_{5},\textit{higher-index terms}_{\geq 5})$ (to $3$ or Pivot State $NR$) $\geq k-c_{n}-1$
$(F_{2},F_{2},\dots,F_{2},\textit{higher-index terms})$ (to Pivot State $NR$) $\geq k-c_{n}-1$
This demonstrates that the number of moves required to traverse from Pivot State NR through an intermediate state and back to a pivot state, thereby reducing the overall tuple length from $k$ to k1k-1, is bounded below by the minimum of the two possible path lengths:
  1. (1)
    Path via $1_{1}$: $(k-c_{n}-2)+(k-c_{n}-1)=2k-2c_{n}-3$
  2. (2)
    Path via $NR_{1}$: $(k-c_{n}-2)+(k-c_{n}-2)=2k-2c_{n}-4$
Thus, the transition from a $k$-tuple to a (k1)(k-1)-tuple starting from the Pivot State $NR$ requires at least $2k-2c_{n}-4$ moves.
Thus, to transition from a $k$-tuple to a (k1)(k-1)-tuple or, equivalently, to traverse between two Pivot States, we require at least $2k-2c_{n}-4$ moves. Summing these bounds yields the total estimate for $M(n)$:
$M(n)$ $~\geq~\frac{n^{2}}{4}+\sum_{k=c_{n}+5}^{\lfloor\frac{n}{2}\rfloor}(2k-2c_{n}-4)$
$~=~\frac{n^{2}}{4}+\lfloor\frac{n}{2}\rfloor\left(\lfloor\frac{n}{2}\rfloor+1\right)-(c_{n}+4)(c_{n}+5)-(2c_{n}+4)\left(\lfloor\frac{n}{2}\rfloor-c_{n}-5\right)$
$~=~\frac{n^{2}}{2}-nc_{n}+O(n)$
@@
We also remark that in the proof above, we can improve a constant before $n\log_{\phi}(n)$ term up to $1/(8\phi)\approx 0.077$ by noting that, asymptotically, $F_{l}\geq n/\phi$ and $l\geq\log_{\phi}(n)$. However, assuming that LGS algorithm indeed gives a maximal game length, numerical simulations suggest that the constant before the $n\log_{\phi}(n)$ term is at least 0.4 and should be much lower that 1, as demonstrated on Figure 5. Hence, while Theorem 1.10 and Theorem 1.11 allow us to establish asymptotic behaviour of $\dfrac{n^{2}}{2}-M(n)$, they contain too crude estimates to give an insight for the explicit constant before $n\log_{\phi}(n)$.
@@
As $n\to\infty$, the distribution of the number of moves in a randomly played game, where all legal moves are equally likely, converges to a log-normal distribution. Under these conditions, the game dynamics are sufficiently symmetric such that each player is equally likely to win.
This conjecture is supported by simulation data presented in Figure 3 and Figure 4, which suggest that the observed move counts exhibit the heavy-tailed, right-skewed characteristics consistent with a log-normal distribution. Furthermore, the symmetry in the outcome frequencies aligns with the theoretical expectation of equal win probabilities for both players under a uniformly random move selection process.
@@ 5. Conclusion
  • Probabilistic Analysis of Random Play: Our empirical evidence suggests a log-normal distribution of move counts under uniform random play. Can this distribution be derived rigorously? What are the deeper probabilistic and ergodic properties of the random dynamics?
@@ 6. Appendix
appears to converge to a number bigger than 0.4 and much lower than 1.